# SwiftRepertoire: Few-Shot Immune-Signature Synthesis via Dynamic Kernel Codes

### v1

@@ 1 Introduction
Early detection of malignancy remains central to improving patient outcomes, and conventional modalities such as organ-specific imaging and circulating biomarkers have shaped current screening practice  Li et al. (2020); Sivapalan et al. (2023); Huang et al. (2024). Recent advances in high-throughput immunogenomics have highlighted the immune repertoire as a complementary window into neoplastic processes because tumor-reactive T cells and their receptors carry signals of antigen exposure and immune selection  Beshnova et al. (2020); Pan and Li (2022); Cai et al. (2024). The variable regions of T cell receptors, and in particular the complementarity determining region three, encode sequence motifs that mediate peptide recognition and can reflect tumor-associated perturbations  Qian et al. (2024); Xu et al. (2022); Sidhom et al. (2021). These biological observations motivate repertoire-based approaches for diagnosis and monitoring that do not depend on tumor somatic markers or limited biomarker panels. Computational methods for repertoire analysis have evolved from statistical enrichment tests and handcrafted feature models to deep learning systems that exploit sequence context and repertoire structure  Smith et al. (2021); Zhang et al. (2024); Cai et al. (2024). Multiple-instance formulations and attention-based pooling have been proposed to aggregate information across repertoires  Kim et al. (2022); Shao et al. (2021); Sidhom et al. (2021). At the same time, protein language models and self-supervised encoders have offered transferable representations that improve downstream binding and classification performance  Dounas et al. (2024); Leem et al. (2022). Despite these advances, several persistent challenges limit practical adoption. First, many repertoire tasks confront severe label sparsity and a long-tail distribution of antigen or disease labels, which causes conventional supervised learners to generalize poorly to rare or unseen tasks  Gao et al. (2023); Wang et al. (2023). Second, adapting large pre-trained encoders to new tasks using full fine-tuning is computationally costly and may overfit when only a few labeled examples exist. Third, parameter-efficient adaptation techniques have been developed, yet there remains a need for methods that can produce immediate, interpretable, and statistically calibrated task-specific behaviors from minimal supervision  Zhao et al. (2023); Lv et al. (2024); Wu et al. (2023). Finally, clinical translation benefits from models that not only predict but also provide motif-level interpretability and calibrated statistical signals for downstream decision making.
@@ 2.1 Repertoire biology and diagnostic motivation
T cell receptor repertoires record antigen-driven selection and contain signals that are clinically useful for cancer detection and immune monitoring. Prior studies report that repertoire-derived summary statistics and clonotype-level features can discriminate cancer from non-cancer cohorts  Beshnova et al. (2020) and reveal spatial or temporal heterogeneity in tumour-infiltrating lymphocytes  Joshi et al. (2019). These biological observations motivate computational pipelines that extract CDR3 motifs and clonotype patterns as complementary diagnostics to imaging and circulating tumour biomarkers  Li et al. (2025); Pan and Li (2022).
@@ 2.2 Representations and deep sequence models
A broad spectrum of sequence-based models has been applied to TCR analysis, including convolutional encoders, transformer and BERT-style architectures, and multiple-instance formulations that pool repertoire evidence. Self-supervised protein language models provide contextual embeddings which, when fine-tuned on downstream tasks, typically yield substantial performance gains  Xu et al. (2022); Zhang et al. (2024). Recent work demonstrates that contrastive, cross-epitope fine-tuning aligns TCR and peptide embeddings in a shared latent space and improves epitope ranking and binding prediction relative to sequence-similarity baselines  Im et al. (2025). Complementary studies show that pretrained PLMs can also reveal confounding biophysical patterns such as viral immune mimicry, highlighting the importance of pairing representational improvements with interpretability and mechanistic analysis  Ofer and Linial (2025); Qian et al. (2024). These advances motivate using fine-tuned protein LMs both for pairwise binding prediction and for repertoire-level classification  Chen et al. (2023).
@@ 2.3 Few-shot learning and task adaptation
The long-tailed distribution of peptide labels and the limited availability of annotated repertoires have driven meta-learning and few-shot approaches for TCR tasks. Memory-augmented meta-learners and task-conditioned modules borrow strength across tasks to generalize to rare or unseen peptides, and pan-peptide meta-learning frameworks have demonstrated robustness to novel antigens in clinically relevant applications  Gao et al. (2023). Other methods emphasize rapid adaptation mechanisms that avoid full fine-tuning while maintaining strong performance from a handful of labeled examples  Wang et al. (2023); Yue et al. (2025).
@@ 2.4 Parameter-efficient adaptation, prototypes, and interpretability
Practical deployment demands methods that trade adaptation capacity against computation and memory. Parameter-efficient adapters, prototype retrieval and fast-weight schemes compress per-task knowledge into compact atoms and enable sparse retrieval to synthesize task-specific parameters with modest overhead  Zhao et al. (2023); Lv et al. (2024). At the same time, interpretable motif discovery and statistically calibrated testing remain essential for translational use; coupling sparse adaptation with motif-level inference yields models that are both predictive and biologically interpretable  Zhang et al. (2022); Wu et al. (2023). Prototype-driven schemes have been combined with motif analysis to provide localized explanations of model decisions  Qian et al. (2024); Pan and Li (2022).
@@ 4.3 Performance comparison on cancer datasets
Lung AAMean Zhang et al. (2024) 0.477 1.000 0.100 0.617 0.805
Lung AAMIL Zhang et al. (2024) 0.739 0.676 0.784 0.685 0.840
Lung BertSingle Zhang et al. (2024) 0.841 0.730 0.922 0.794 0.935
Lung MethPriorGCN Ni et al. (2025) 0.810 0.650 0.900 0.500 0.800
Lung BertTCR Zhang et al. (2024) 0.898 0.784 0.980 0.866 0.959
@@
THCA AAMean Zhang et al. (2024) 0.500 0.975 0.087 0.645 0.693
THCA AAMIL Zhang et al. (2024) 0.802 0.625 0.957 0.746 0.899
THCA BertSingle Zhang et al. (2024) 0.860 0.725 0.978 0.829 0.979
THCA BertTCR Zhang et al. (2024) 0.930 0.875 0.978 0.921 0.990
THCA MethPriorGCN Ni et al. (2025) 0.8060 0.600 0.900 0.4787 0.7993
@@ 4.4 Comparative analysis
Lung DeepCAT Beshnova et al. (2020) 0.620 0.871 0.388 0.688 0.720
Lung DeepLION Xu et al. (2022) 0.620 0.887 0.373 0.806 0.762
Lung BertSingle Zhang et al. (2024) 0.798 0.871 0.731 0.806 0.907
Lung BertTCR Zhang et al. (2024) 0.899 0.952 0.851 0.901 0.972
Lung DeepLION2 Qian et al. (2024) 0.809 0.736 0.865 0.770 0.880
Lung DeepTCR Sidhom et al. (2021) 0.721 0.393 0.968 0.617 0.836
Lung MINN_SA Kim et al. (2022) 0.655 0.620 0.711 0.685 0.788
Lung TransMIL Shao et al. (2021) 0.757 0.687 0.814 0.746 0.820
Lung BiFormer Zhu et al. (2023) 0.768 0.716 0.812 0.746 0.836
@@
THCA DeepCAT Beshnova et al. (2020) 0.600 0.917 0.308 0.687 0.731
THCA DeepLION Xu et al. (2022) 0.680 0.583 0.769 0.636 0.769
THCA BertSingle Zhang et al. (2024) 0.720 0.583 0.846 0.667 0.785
THCA BertTCR Zhang et al. (2024) 0.900 0.875 0.923 0.894 0.912
THCA DeepTCR Sidhom et al. (2021) 0.733 0.481 0.891 0.617 0.860
THCA MINN_SA Kim et al. (2022) 0.740 0.542 0.874 0.608 0.843
THCA TransMIL Shao et al. (2021) 0.816 0.704 0.887 0.746 0.888
THCA BiFormer Zhu et al. (2023) 0.840 0.729 0.911 0.770 0.917
THCA DeepLION2 Qian et al. (2024) 0.886 0.751 0.973 0.908 0.933

### v2

@@ 1 Introduction
Early detection of malignancy remains central to improving patient outcomes, and conventional modalities such as organ-specific imaging and circulating biomarkers have shaped current screening practice  [11, 20, 6]. Recent advances in high-throughput immunogenomics have highlighted the immune repertoire as a complementary window into neoplastic processes because tumor-reactive T cells and their receptors carry signals of antigen exposure and immune selection  [1, 16, 2]. The variable regions of T cell receptors, and in particular the complementarity determining region three, encode sequence motifs that mediate peptide recognition and can reflect tumor-associated perturbations  [17, 24, 19]. These biological observations motivate repertoire-based approaches for diagnosis and monitoring that do not depend on tumor somatic markers or limited biomarker panels. Computational methods for repertoire analysis have evolved from statistical enrichment tests and handcrafted feature models to deep learning systems that exploit sequence context and repertoire structure  [21, 26, 2]. Multiple-instance formulations and attention-based pooling have been proposed to aggregate information across repertoires  [9, 18, 19]. At the same time, protein language models and self-supervised encoders have offered transferable representations that improve downstream binding and classification performance  [4, 10]. Despite these advances, several persistent challenges limit practical adoption. First, many repertoire tasks confront severe label sparsity and a long-tail distribution of antigen or disease labels, which causes conventional supervised learners to generalize poorly to rare or unseen tasks  [5, 22]. Second, adapting large pre-trained encoders to new tasks using full fine-tuning is computationally costly and may overfit when only a few labeled examples exist. Third, parameter-efficient adaptation techniques have been developed, yet there remains a need for methods that can produce immediate, interpretable, and statistically calibrated task-specific behaviors from minimal supervision  [28, 13, 23]. Finally, clinical translation benefits from models that not only predict but also provide motif-level interpretability and calibrated statistical signals for downstream decision making.
@@ 2.1 Repertoire biology and diagnostic motivation
T cell receptor repertoires record antigen-driven selection and contain signals that are clinically useful for cancer detection and immune monitoring. Prior studies report that repertoire-derived summary statistics and clonotype-level features can discriminate cancer from non-cancer cohorts  [1] and reveal spatial or temporal heterogeneity in tumour-infiltrating lymphocytes  [8]. These biological observations motivate computational pipelines that extract CDR3 motifs and clonotype patterns as complementary diagnostics to imaging and circulating tumour biomarkers  [12, 16].
@@ 2.2 Representations and deep sequence models
A broad spectrum of sequence-based models has been applied to TCR analysis, including convolutional encoders, transformer and BERT-style architectures, and multiple-instance formulations that pool repertoire evidence. Self-supervised protein language models provide contextual embeddings which, when fine-tuned on downstream tasks, typically yield substantial performance gains  [24, 26]. Recent work demonstrates that contrastive, cross-epitope fine-tuning aligns TCR and peptide embeddings in a shared latent space and improves epitope ranking and binding prediction relative to sequence-similarity baselines  [7]. Complementary studies show that pretrained PLMs can also reveal confounding biophysical patterns such as viral immune mimicry, highlighting the importance of pairing representational improvements with interpretability and mechanistic analysis  [15, 17]. These advances motivate using fine-tuned protein LMs both for pairwise binding prediction and for repertoire-level classification  [3].
@@ 2.3 Few-shot learning and task adaptation
The long-tailed distribution of peptide labels and the limited availability of annotated repertoires have driven meta-learning and few-shot approaches for TCR tasks. Memory-augmented meta-learners and task-conditioned modules borrow strength across tasks to generalize to rare or unseen peptides, and pan-peptide meta-learning frameworks have demonstrated robustness to novel antigens in clinically relevant applications  [5]. Other methods emphasize rapid adaptation mechanisms that avoid full fine-tuning while maintaining strong performance from a handful of labeled examples  [22, 25].
@@ 2.4 Parameter-efficient adaptation, prototypes, and interpretability
Practical deployment demands methods that trade adaptation capacity against computation and memory. Parameter-efficient adapters, prototype retrieval and fast-weight schemes compress per-task knowledge into compact atoms and enable sparse retrieval to synthesize task-specific parameters with modest overhead  [28, 13]. At the same time, interpretable motif discovery and statistically calibrated testing remain essential for translational use; coupling sparse adaptation with motif-level inference yields models that are both predictive and biologically interpretable  [27, 23]. Prototype-driven schemes have been combined with motif analysis to provide localized explanations of model decisions  [17, 16].
@@ 4.3 Performance comparison on cancer datasets
Lung AAMean [26] 0.477 1.000 0.100 0.617 0.805
Lung AAMIL [26] 0.739 0.676 0.784 0.685 0.840
Lung BertSingle [26] 0.841 0.730 0.922 0.794 0.935
Lung MethPriorGCN [14] 0.810 0.650 0.900 0.500 0.800
Lung BertTCR [26] 0.898 0.784 0.980 0.866 0.959
@@
THCA AAMean [26] 0.500 0.975 0.087 0.645 0.693
THCA AAMIL [26] 0.802 0.625 0.957 0.746 0.899
THCA BertSingle [26] 0.860 0.725 0.978 0.829 0.979
THCA BertTCR [26] 0.930 0.875 0.978 0.921 0.990
THCA MethPriorGCN [14] 0.8060 0.600 0.900 0.4787 0.7993
@@ 4.4 Comparative analysis
Lung DeepCAT [1] 0.620 0.871 0.388 0.688 0.720
Lung DeepLION [24] 0.620 0.887 0.373 0.806 0.762
Lung BertSingle [26] 0.798 0.871 0.731 0.806 0.907
Lung BertTCR [26] 0.899 0.952 0.851 0.901 0.972
Lung DeepLION2 [17] 0.809 0.736 0.865 0.770 0.880
Lung DeepTCR [19] 0.721 0.393 0.968 0.617 0.836
Lung MINN_SA [9] 0.655 0.620 0.711 0.685 0.788
Lung TransMIL [18] 0.757 0.687 0.814 0.746 0.820
Lung BiFormer [29] 0.768 0.716 0.812 0.746 0.836
@@
THCA DeepCAT [1] 0.600 0.917 0.308 0.687 0.731
THCA DeepLION [24] 0.680 0.583 0.769 0.636 0.769
THCA BertSingle [26] 0.720 0.583 0.846 0.667 0.785
THCA BertTCR [26] 0.900 0.875 0.923 0.894 0.912
THCA DeepTCR [19] 0.733 0.481 0.891 0.617 0.860
THCA MINN_SA [9] 0.740 0.542 0.874 0.608 0.843
THCA TransMIL [18] 0.816 0.704 0.887 0.746 0.888
THCA BiFormer [29] 0.840 0.729 0.911 0.770 0.917
THCA DeepLION2 [17] 0.886 0.751 0.973 0.908 0.933